OpenStudy (anonymous):

in the game of yahtzee,, on the first roll five dice are tossed simultaneously. what is the probability of rolling five of a kind on the first roll?

OpenStudy (dumbcow):

the probability of rolling any given number is 1/6. there are 5 dice so multiply the probabilities P = (1/6)^5 Multiply by 6 because there are 6 possible ways of rolling 5 of a kind P = 6*(1/6)^5 = (1/6)^4

OpenStudy (anonymous):

thank you very much :)